OpenStudy (anonymous):

Which of the following is the product of (a + b)(a – b)? a. a^2+b^2 b. a^2b^2 c. a^2+2ab+b^2 d. a^2-2ab+b^2

OpenStudy (solomonzelman):

\(\large\color{ midnightblue }{ \rm (a+b)(a-b) =a^2-b^2 }\)

OpenStudy (solomonzelman):

\(\large\color{blue }{ \rm (a+b)(a-b) =a^2-b^2 }\) Also, \(\large\color{blue }{ \rm (a+b)(a+b) =a^2+2ab+b^2 }\) \(\large\color{blue }{ \rm (a-b)(a-b) =a^2-2ab+b^2 }\) \(\large\color{blue }{ \rm (a+i~b~)(~a-i~b) =a^2+b^2 }\)

OpenStudy (anonymous):

k